### v2

# Chiral non-Abelian vortex molecules in dense QCD

Abstract
The color-flavor locked phase of high density QCD admits non-Abelian vortices, that are superfluid vortices with confined color-magnetic fluxes. Vortex solutions were thus far constructed in an axially symmetric Ansatz with common vortex winding in the left- and right-handed diquark condensates. In this paper, we explore vortex solutions without any Ansatz. In addition to axially symmetric configurations known before, we find that the axial symmetry is broken in certain parameter regions; in one case a single vortex is split into two chiral non-Abelian vortices, i.e. vortices with winding only in the left or right-handed diquark condensate, and they are connected by one or two domain walls forming a non-Abelian vortex molecule. In the other case, a chiral non-Abelian vortex molecule is confined inside a domain wall elongated to spatial infinity.
1 Introduction
Exploring states of matter in extreme conditions is an important problem in modern physics. In particular, nuclear and quark matter in extreme conditions such as high density, strong magnetic field, and rapid rotation are considered to be pivotal for a full understanding of the interior of compact stars such as neutron stars. The quark matter described by quantum chromodynamics (QCD) at high densities and low temperatures is expected to exhibit color superconductivity as well as superfluidity Alford:2007xm . In the limit of extremely high density, a phase with three-flavor symmetric matter called the color-flavor locked (CFL) phase is realized Alford:1998mk , while the two-flavor superconducting (2SC) phase Alford:1997zt ; Rapp:1997zu is also proposed for the intermediate-density region of the QCD phase diagram. In metallic superconductors and superfluids, quantum vortices or magnetic flux tubes play significant roles in color-superconducting quark matter, as reviewed in ref. Eto:2013hoa . In the CFL phase, the most fundamental stable vortices are non-Abelian vortices carrying color-magnetic fluxes and 1/3 superfluid circulation Eto:2013hoa ; Balachandran:2005ev ; Nakano:2007dr ; Nakano:2008dc ; Eto:2009kg ; Eto:2009bh ; Eto:2009tr .111 Similar non-Abelian vortices were studied in supersymmetric QCD Hanany:2003hp ; Auzzi:2003fs ; Hanany:2004ea ; Shifman:2004dr ; Eto:2004rz ; Eto:2005yh (see refs. Tong:2005un ; Eto:2006pg ; Shifman:2007ce ; Shifman:2009zz for a review) and in two-Higgs doublet models (as an extension of the Standard Model) Eto:2018hhg ; Eto:2018tnk ; Eto:2019hhf ; Eto:2020hjb ; Eto:2020opf ; Eto:2021dca . They are different from the system studied here, in the sense that the overall $\operatorname{U}(1)$ symmetry is gauged in supersymmetric QCD, and it is an approximate symmetry in two-Higgs doublet models. Superfluid vortices without fluxes called Abelian vortices Forbes:2001gj ; Iida:2002ev are dynamically unstable to decay into three non-Abelian vortices Nakano:2007dr ; Cipriani:2012hr ; Alford:2016dco . A non-Abelian vortex confines bosonic and fermionic gapless modes in its core; the former are Nambu-Goldstone ${\mathbb{C}}P^{2}(\simeq{\rm SU}(3)/[{\rm SU}(2)\times{\rm U}(1)])$ modes originated from spontaneous breaking of the CFL symmetry in the vicinity of the core Nakano:2007dr ; Eto:2009bh ; Eto:2013hoa ; Eto:2009tr while the latter are three Majorana fermions Yasui:2010yw ; Fujiwara:2011za ; Chatterjee:2016ykq . Under a rapid rotation, there appear a huge number of vortices (about $10^{19}$ for typical neutron stars) aligned along the rotation axis, forming a vortex lattice with the lattice spacing of the order of a micrometer Kobayashi:2013axa ; Hirono:2012ki . Such vortices penetrate via crossover between the CFL phase and hyperon nuclear matter within a quark-hadron continuity Cipriani:2012hr ; Alford:2018mqj ; Chatterjee:2018nxe ; Chatterjee:2019tbz ; Cherman:2018jir ; Hirono:2018fjr ; Hirono:2019oup ; Cherman:2020hbe without Alford:2018mqj or with Boojums Cipriani:2012hr ; Chatterjee:2018nxe ; Chatterjee:2019tbz . This direction has stimulated further studies of the Higgs-confinement continuity in the presence of a vortex Cherman:2020hbe ; Hayashi:2023sas ; Cherman:2024exo ; Hayata:2024nrl .
The CFL phase is characterized by the two diquark condensates of left and right-handed quarks $q_{\rm L,R}$, $(\Phi_{\rm L,R})_{\alpha a}\sim\epsilon_{\alpha\beta\gamma}\epsilon_{abc}q_{\rm L,R}^{\beta b}q_{\rm L,R}^{\gamma c}$ with the color indices $\alpha,\beta,\gamma=r,g,b$ and the flavor indices $a,b,c=u,d,s$. In the ground state, they simultaneously develop vacuum expectation values (VEVs) as $\Phi_{\rm L}=-\Phi_{\rm R}$. In the previous studies Balachandran:2005ev ; Eto:2009kg ; Eto:2013hoa , non-Abelian vortices were constructed in a simplified setup with the assumption $\Phi_{\rm L}=-\Phi_{\rm R}$ in the entire region, including vortex cores and the axisymmetry around the vortices. In order to investigate more general vortex configurations, it is convenient to recall a simpler setup of condensed matter systems with simultaneous condensation of two components $\Phi_{1}$ and $\Phi_{2}$. Examples contain two-gap superconductors Babaev:2001hv ; doi:10.1143/JPSJ.70.2844 ; PhysRevLett.88.017002 ; Goryo_2007 , chiral $P$-wave superconductors PhysRevLett.80.5184 ; PhysRevB.86.060514 and two-component Bose-Einstein condensates (BECs) Son:2001td ; Kasamatsu:2004tvg ; Cipriani:2013nya ; Tylutki:2016mgy ; Eto:2017rfr ; Eto:2019uhe ; Kobayashi:2018ezm , in which the overall phase is a gauge (global) symmetry in the superconductors (BECs), and the relative phase is an approximate global symmetry explicitly broken by a Josephson term (Rabi coupling) $\Phi_{1}^{*}\Phi_{2}+{\rm c.c.}$ (or $\Phi_{1}^{*2}\Phi_{2}^{2}+{\rm c.c.}$ for chiral $P$-wave superconductors PhysRevLett.80.5184 ; PhysRevB.86.060514 ). A singly-quantized vortex has winding in both components: $\Phi_{1}=\Phi_{2}\sim e^{\mathrm{i}\theta}$ at the same position with the azimuthal angle $\theta$. Depending on the situation, it can be split into two half-quantized vortices $(\Phi_{1},\Phi_{2})\sim(e^{\mathrm{i}\theta},1)$ and $(\Phi_{1},\Phi_{2})\sim(1,e^{\mathrm{i}\theta})$ at different positions connected by a sine-Gordon soliton doi:10.1143/JPSJ.70.2844 ; PhysRevLett.88.017002 ; Goryo_2007 ; Son:2001td (or two domain walls for chiral $P$-wave superconductors PhysRevB.86.060514 ) forming a vortex molecule.222 See ref. Chatterjee:2019jez for the case of more general higher-order Josephson-like terms and refs. Eto:2012rc ; Eto:2013spa for the cases of more components and corresponding vortex-molecule structures.
In this paper, we explore non-Abelian vortex solutions without imposing an Ansatz; no assumption of $\Phi_{\rm L}=-\Phi_{\rm R}$ nor of axisymmetry, unlike the previous studies Balachandran:2005ev ; Eto:2009kg ; Eto:2013hoa . We find that for certain parameter choices, a single non-Abelian vortex is split into two vortex cores breaking the axisymmetry, each of which has a winding only in left $\Phi_{\rm L}$ or in the right $\Phi_{\rm R}$ condensates. Such constituents are called chiral non-Abelian vortices because a condensate of only left or right chirality has a vortex-winding Eto:2021nle . Each chiral non-Abelian vortex is attached by one or two domain walls Eto:2013hoa ; Eto:2013bxa , due to the fact that axial and chiral symmetries are explicitly broken by the mass and axial-anomaly terms. The two chiral non-Abelian vortices with opposite chiralities are connected by one or two domain walls forming a non-Abelian vortex molecule. In certain parameter regions, the energy remains axisymmetric, but the vortex centers are not coincident – the axial symmetry is hence broken. We also find, in another parameter region, that the two chiral non-Abelian vortices are confined on a single domain wall, with one domain wall connecting them and two domain walls elongated to spatial infinities.
This paper is organized as follows. In sec. 2 we summarize the Ginzburg-Landau (GL) energy functional of the CFL phase of dense QCD and the ground states of the CFL phase. In sec. 3 we investigate vortex solutions without imposing an Ansatz, and find a non-Abelian vortex molecule in which two chiral vortices of opposite chiralities are connected by one or two domain walls, as well as a vortex molecule confined inside a single domain wall. Sec. 4 is devoted to a summary and discussion.
2 Color-flavor-locked phase of 3-flavor dense QCD
In this section, we review the GL model for the CFL phase to fix our notations, and summarize the ground state structures.
2.1 Ginzburg-Landau model for the color-flavor locked phase
Let $q_{{\rm L},{\rm R}}$ be left- and right-handed quarks. Then, the CFL phase is characterized by the simultaneous diquark condensates of the left and right chiralities:
$(\Phi_{{\rm L},{\rm R}})_{\alpha a}\sim\epsilon_{\alpha\beta\gamma}\epsilon_{abc}q_{{\rm L},{\rm R}}^{\beta b}q_{{\rm L},{\rm R}}^{\gamma c},$ (1)
which are 3-by-3 matrices of scalar fields with $\alpha,\beta,\gamma=r,g,b=1,2,3$ being color indices and $a,b,c=u,d,s=1,2,3$ being flavor indices. The ${\rm SU}(3)_{\rm C}$ color symmetry and ${\rm U}(1)_{\rm B}$ baryon symmetry are exact while the ${\rm SU}(3)_{\rm L}\times{\rm SU}(3)_{\rm R}$ chiral symmetry and ${\rm U}(1)_{\rm A}$ axial symmetry are approximate. These symmetries act on the condensates as
$\Phi_{\rm L}\to e^{i\theta_{\rm B}+i\theta_{\rm A}}g_{\rm C}\Phi_{\rm L}U_{\rm L}^{\dagger},\quad\Phi_{\rm R}\to e^{i\theta_{\rm B}-i\theta_{\rm A}}g_{\rm C}\Phi_{\rm R}U_{\rm R}^{\dagger}$
$g_{\rm C}\in{\rm SU}(3)_{\rm C},\quad U_{\rm L,R}\in{\rm SU}(3)_{\rm L,R},\quad e^{i\theta_{\rm B}}\in{\rm U}(1)_{\rm B},\quad e^{i\theta_{\rm A}}\in{\rm U}(1)_{\rm A}.$ (2)
The vector symmetry ${\rm SU}(3)_{\rm L+R}$ defined by the condition $U_{\rm L}=U_{\rm R}$ is a subgroup of the chiral symmetry ${\rm SU}(3)_{\rm L}\times{\rm SU}(3)_{\rm R}$, and the rest of the generators parametrize Nambu-Goldstone bosons for the chiral symmetry breaking as the coset space
$[{\rm SU}(3)_{\rm L}\times{\rm SU}(3)_{\rm R}]/{\rm SU}(3)_{\rm L+R}\simeq{\rm SU}(3).$
The static Hamiltonian (energy functional) of the GL model takes the form [^3] The $\lambda_{2}$ and $\lambda_{3}$ terms are divided by $N=3$ for convenience, compared e.g. with ref. Eto:2021nle .
$E$ $=\frac{1}{2g^{2}}|\mkern-1.0mu|F|\mkern-1.0mu|^{2}+|\mkern-1.0mu|{\mathrm{d}}_{A}\Phi_{L}|\mkern-1.0mu|^{2}+|\mkern-1.0mu|{\mathrm{d}}_{A}\Phi_{R}|\mkern-1.0mu|^{2}+V,$ (3)
$V$ $=-\frac{m^{2}}{2}\int_{M}\star\operatorname{Tr}\big{(}\Phi_{\rm L}^{\dagger}\Phi_{\rm L}+\Phi_{\rm R}^{\dagger}\Phi_{\rm R}\big{)}+\frac{\lambda_{1}}{4}\int_{M}\star\operatorname{Tr}\left[\big{(}\Phi_{\rm L}^{\dagger}\Phi_{\rm L}\big{)}^{2}+\big{(}\Phi_{\rm R}^{\dagger}\Phi_{\rm R}\big{)}^{2}\right]$
$\phantom{=\ }+\frac{\lambda_{2}}{12}\int_{M}\star\left(\left(\operatorname{Tr}\big{[}\Phi_{\rm L}^{\dagger}\Phi_{L}\big{]}\right)^{2}+\left(\operatorname{Tr}\big{[}\Phi_{\rm R}^{\dagger}\Phi_{\rm R}\big{]}\right)^{2}\right)+\frac{\lambda_{3}}{6}\int_{M}\star\operatorname{Tr}\big{[}\Phi_{\rm L}^{\dagger}\Phi_{\rm L}\big{]}\operatorname{Tr}\big{[}\Phi_{\rm R}^{\dagger}\Phi_{\rm R}\big{]}$
$\phantom{=\ }+\frac{\lambda_{4}}{2}\int_{M}\star\operatorname{Tr}\big{[}\Phi_{\rm R}^{\dagger}\Phi_{\rm L}\Phi_{\rm L}^{\dagger}\Phi_{\rm R}\big{]}+\gamma_{1}\int_{M}\star\operatorname{Tr}\big{(}\Phi_{\rm L}^{\dagger}\Phi_{\rm R}+\Phi_{\rm L}\Phi_{\rm R}^{\dagger}\big{)}$
$\phantom{=\ }+\gamma_{2}\int_{M}\star\operatorname{Tr}\left[\big{(}\Phi_{\rm L}^{\dagger}\Phi_{\rm R}\big{)}^{2}+\big{(}\Phi_{\rm R}^{\dagger}\Phi_{\rm L}\big{)}^{2}\right]+\gamma_{3}\int_{M}\star\left(\det\big{(}\Phi_{\rm L}^{\dagger}\Phi_{\rm R}\big{)}+\det\big{(}\Phi_{\rm R}^{\dagger}\Phi_{\rm L}\big{)}\right),$ (4)
where the inner product on $M=\mathbb{R}^{2}$ (the plane orthogonal to vortices) is defined as
$\langle X,Y\rangle:=\operatorname{Tr}\int_{M}X^{\dagger}\wedge\star Y,$ (5)
where $X,Y$ are both $r$-forms as well as 3-by-3 complex matrices and $\star$ is the Hodge star mapping $r$-forms to ($2-r$)-forms with the property $\star\star=(-1)^{r}$ (in two dimensions). The norm squared is then defined as
$|\mkern-1.0mu|X|\mkern-1.0mu|^{2}:=\langle X,X\rangle,$ (6)
and the integral of the Hodge dual of a scalar is simply the normal integral with the volume form: $\int_{M}\star Z=\int_{M}Z\,{\mathrm{d}}^{2}x$ with $Z$ a 0-form. The field-strength 2-form for the SU(3)C color gauge field is defined as
$F={\mathrm{d}}A-A\wedge A=\frac{1}{2}F_{ij}{\mathrm{d}}x^{ij},$ (7)
with the short-hand notation ${\mathrm{d}}x^{ij}:={\mathrm{d}}x^{i}\wedge{\mathrm{d}}x^{j}$, $i,j=1,2$ and the gauge covariant derivative
${\mathrm{d}}_{A}\Phi_{{\rm L},{\rm R}}:=d\Phi_{{\rm L},{\rm R}}-A\Phi_{{\rm L},{\rm R}},$ (8)
where $A=A_{i}{\mathrm{d}}x^{i}$ is an anti-Hermitian 1-form, i.e. $A^{\dagger}=-A$ and is $\mathfrak{su}(3)$-valued. This implies that it is also traceless.
For convenience, we provide a few expressions in component form
$\frac{1}{2g^{2}}|\mkern-1.0mu|F|\mkern-1.0mu|^{2}$ $=-\frac{1}{4g^{2}}\int_{M}\operatorname{Tr}\left[F_{ij}F^{ij}\right]{\mathrm{d}}^{2}x\geq 0,$ (9)
$|\mkern-1.0mu|{\mathrm{d}}_{A}\Phi_{\rm L}|\mkern-1.0mu|^{2}$ $=\int_{M}\operatorname{Tr}\left[(\partial_{i}\Phi_{\rm L}-A_{i}\Phi_{\rm L})^{\dagger}(\partial^{i}\Phi_{\rm L}-A^{i}\Phi_{\rm L})\right]{\mathrm{d}}^{2}x,$ (10)
where spatial indices $i,j$ are lowered (raised) by the (inverse) metric $g_{ij}=\delta_{ij}$ ($g^{ij}=\delta^{ij}$). Note the negative sign in front of $F_{ij}F^{ij}$ is due to the anti-Hermitian gauge field, i.e. $F^{\dagger}=-F$.
The GL parameters in the GL model in eq. (4) were microscopically calculated in the asymptotically high-density region of QCD Iida:2000ha ; Iida:2001pg ; Giannakis:2001wz . In this paper, we leave those parameters as free GL parameters.
2.2 First variation
In order to obtain the equations of motion, we perform a first variation of the energy functional. Let $(A_{\tau},\Phi_{{\rm L},\tau},\Phi_{{\rm R},\tau})$ be smooth variations of the fields $(A,\Phi_{\rm L},\Phi_{\rm R})$ for all $\tau$. Denote by $\alpha=\partial_{\tau}A_{\tau}|_{\tau=0}$, $\beta_{\rm L}=\partial_{\tau}\Phi_{{\rm L},\tau}|_{\tau=0}$ and $\beta_{\rm R}=\partial_{\tau}\Phi_{{\rm R},\tau}|_{\tau=0}$. We thus obtain
$\frac{{\mathrm{d}}}{{\mathrm{d}}\tau}\bigg{|}_{\tau=0}\!E$ $=\langle\beta_{\rm L},\operatorname{eom}_{\Phi_{\rm L}}\rangle_{L^{2}(M)}+\langle\operatorname{eom}_{\Phi_{\rm L}},\beta_{\rm L}\rangle_{L^{2}(M)}+\langle\beta_{\rm R},\operatorname{eom}_{\Phi_{\rm R}}\rangle_{L^{2}(M)}+\langle\operatorname{eom}_{\Phi_{\rm R}},\beta_{\rm R}\rangle_{L^{2}(M)}$
$\phantom{=\ }+\langle\alpha,\operatorname{eom}_{A}\rangle_{L^{2}(M)}-\frac{1}{g^{2}}\operatorname{Tr}\int_{M}{\mathrm{d}}\left(\alpha\wedge\star F\right)$
$\phantom{=\ }+\operatorname{Tr}\int_{M}{\mathrm{d}}\left(\beta_{\rm L}^{\dagger}\star{\mathrm{d}}_{A}\Phi_{\rm L}+\star{\mathrm{d}}_{A}\Phi_{\rm L}^{\dagger}\beta_{\rm L}+\beta_{\rm R}^{\dagger}\star{\mathrm{d}}_{A}\Phi_{\rm R}+\star{\mathrm{d}}_{A}\Phi_{\rm R}^{\dagger}\beta_{\rm R}\right),$ (11)
with the equations of motion
$\operatorname{eom}_{\Phi_{\rm L}}$ $=\delta_{A}{\mathrm{d}}_{A}\Phi_{\rm L}-\frac{m^{2}}{2}\Phi_{\rm L}+\frac{\lambda_{1}}{2}\Phi_{\rm L}\Phi_{\rm L}^{\dagger}\Phi_{\rm L}+\frac{\lambda_{2}}{6}\Phi_{\rm L}\operatorname{Tr}(\Phi_{\rm L}^{\dagger}\Phi_{\rm L})+\frac{\lambda_{3}}{6}\Phi_{\rm L}\operatorname{Tr}(\Phi_{\rm R}^{\dagger}\Phi_{\rm R})$
$\phantom{=\ }+\frac{\lambda_{4}}{2}\Phi_{\rm R}\Phi_{\rm R}^{\dagger}\Phi_{\rm L}+\gamma_{1}\Phi_{\rm R}+2\gamma_{2}\Phi_{\rm R}\Phi_{\rm L}^{\dagger}\Phi_{\rm R}+\gamma_{3}\Xi_{\rm L},$ (12)
$\operatorname{eom}_{\Phi_{\rm R}}$ $=\delta_{A}{\mathrm{d}}_{A}\Phi_{\rm R}-\frac{m^{2}}{2}\Phi_{\rm R}+\frac{\lambda_{1}}{2}\Phi_{\rm R}\Phi_{\rm R}^{\dagger}\Phi_{\rm R}+\frac{\lambda_{2}}{6}\Phi_{\rm R}\operatorname{Tr}(\Phi_{\rm R}^{\dagger}\Phi_{\rm R})+\frac{\lambda_{3}}{6}\Phi_{\rm R}\operatorname{Tr}(\Phi_{\rm L}^{\dagger}\Phi_{\rm L})$
$\phantom{=\ }+\frac{\lambda_{4}}{2}\Phi_{\rm L}\Phi_{\rm L}^{\dagger}\Phi_{\rm R}+\gamma_{1}\Phi_{\rm L}+2\gamma_{2}\Phi_{\rm L}\Phi_{\rm R}^{\dagger}\Phi_{\rm L}+\gamma_{3}\Xi_{\rm R},$ (13)
$\operatorname{eom}_{A}$ $=\frac{1}{g^{2}}\left[\delta F+\star A\wedge\star F-\star(\star F\wedge A)\right]-{\mathrm{d}}_{A}\Phi_{\rm L}\Phi_{\rm L}^{\dagger}+\Phi_{\rm L}{\mathrm{d}}_{A}\Phi_{\rm L}^{\dagger}-{\mathrm{d}}_{A}\Phi_{\rm R}\Phi_{\rm R}^{\dagger}+\Phi_{\rm R}{\mathrm{d}}_{A}\Phi_{\rm R}^{\dagger}$
$=\bigg{(}-\frac{1}{g^{2}}\partial^{i}F_{ij}+\frac{1}{g^{2}}[A^{i},F_{ij}]-(\partial_{j}\Phi_{\rm L}-A_{j}\Phi_{L})\Phi_{\rm L}^{\dagger}+\Phi_{\rm L}(\partial_{j}\Phi_{\rm L}^{\dagger}+\Phi_{\rm L}^{\dagger}A_{j})$
[…]
which are two 0-forms and a 1-form that vanish when the equations of motion are satisfied. In eq. (11), the last two terms, being total derivatives, vanish by the assumption of the smooth variations $\alpha$ and $\beta$ having compact support. [^4] From a physical point of view, the finite-energy condition requires $F$ and ${\mathrm{d}}_{A}\Phi_{{\rm L},{\rm R}}$ to vanish at spatial infinity and hence the total derivatives vanish. $\delta$ is the coderivative and $\delta_{A}$ is the gauge covariant coderivative; writing out the $\delta_{A}{\mathrm{d}}_{A}$ we obtain
$\delta_{A}{\mathrm{d}}_{A}\Phi_{\rm L}$ $=-\star{\mathrm{d}}_{A}\star{\mathrm{d}}_{A}\Phi_{\rm L}=-(\partial_{i}-A_{i})(\partial^{i}-A^{i})\Phi_{\rm L}.$ (15)
Finally, the matrices $\Xi_{{\rm L},{\rm R}}$ which are the variation of the determinant, can be written as
$(\Xi_{\rm L})_{\alpha a}$ $=\frac{1}{2}\epsilon_{abc}\epsilon_{def}(\Phi_{\rm R})_{\alpha d}(\Phi_{\rm L}^{\dagger}\Phi_{R})_{be}(\Phi_{\rm L}^{\dagger}\Phi_{\rm R})_{cf},$ (16)
$(\Xi_{\rm R})_{\alpha a}$ $=\frac{1}{2}\epsilon_{abc}\epsilon_{def}(\Phi_{\rm L})_{\alpha d}(\Phi_{\rm R}^{\dagger}\Phi_{L})_{be}(\Phi_{\rm R}^{\dagger}\Phi_{\rm L})_{cf}.$ (17)
Having the equations of motion in hand, we can read off the perturbative masses from the linearized equations:
$m_{\Phi}=\frac{m}{\sqrt{2}},\qquad m_{A}=g\sqrt{2(v_{\rm L}^{2}+v_{\rm R}^{2})},$ (18)
with the two VEVs $v_{\rm L}=\langle\Phi_{\rm L}\rangle$ and $v_{\rm R}=\langle\Phi_{\rm R}\rangle$ being determined by the ground state equations, see below.
2.3 Ground states
Let us assume that the ground states are given by diagonal matrices, which thus do not break the $\operatorname{SU}(3)_{\rm C+L+R}$ symmetry. We will further use the $\operatorname{U}(1)_{\rm A,B}$ symmetries to set the phases of the VEVs to zero (or $\pi$) if possible.
Let us consider the ground state with only $m\neq 0$, $\lambda_{1}>0$, $\lambda_{2}>0$. The ground state solution compatible with this condition is unique
$\Phi_{\rm L}=-\Phi_{\rm R}=v\mathbf{1}_{3},\qquad v=\frac{m}{\sqrt{\lambda_{1}+\lambda_{2}}},\qquad\langle V\rangle=-\frac{3m^{4}}{2(\lambda_{1}+\lambda_{2})},$ (19)
where $\langle V\rangle$ is the potential value at the VEV. The symmetry is broken to the CFL symmetry ${\rm SU}(3)_{\rm C+L+R}$ given by $g_{\rm C}=U_{\rm L}=U_{\rm R}$ in eq. (2).
Turning on the mixed Hermitian terms, $\lambda_{3}>0$ and $\lambda_{4}>0$, two competing ground states appear:
$\Phi_{\rm L}=-\Phi_{\rm R}=u\mathbf{1}_{3},\quad u=\frac{m}{\sqrt{\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}}},\quad\langle V\rangle=-\frac{3m^{4}}{2(\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4})},$ (20)
and
$\left\{\begin{array}[]{l}\Phi_{\rm L}=v\mathbf{1}_{3}\\ \Phi_{\rm R}=\mathbf{0}_{3}\end{array}\right\}\ \ \textrm{or}\ \ \left\{\begin{array}[]{l}\Phi_{\rm L}=\mathbf{0}_{3}\\ \Phi_{\rm R}=v\mathbf{1}_{3}\end{array}\right\},\quad v=\frac{m}{\sqrt{\lambda_{1}+\lambda_{2}}},\quad\langle V\rangle=-\frac{3m^{4}}{4(\lambda_{1}+\lambda_{2})}.$ (25)
The condition for the ground state being either the $v$ or the $u$ ground state is
$\lambda_{1}+\lambda_{2}>\lambda_{3}+\lambda_{4}$ $:\qquad\Rightarrow\qquad\Phi_{\rm L}=-\Phi_{\rm R}=u\mathbf{1}_{3},$ (26)
$\lambda_{1}+\lambda_{2}<\lambda_{3}+\lambda_{4}$ :$:\qquad\Rightarrow\qquad\left\{\begin{array}[]{l}\Phi_{\rm L}=v\mathbf{1}_{3}\\ \Phi_{\rm R}=\mathbf{0}_{3}\end{array}\right\}\quad\textrm{or}\quad\left\{\begin{array}[]{l}\Phi_{\rm L}=\mathbf{0}_{3}\\ \Phi_{\rm R}=v\mathbf{1}_{3}\end{array}\right\}.$ (31)
The former is still the conventional CFL ground state with the CFL symmetry, while the latter ground state of the $v$-ground state (requiring large $\lambda_{3}$ or $\lambda_{4}$) contains two degenerate ground states and hence a domain wall that connects them.
We will now turn on $\gamma_{1}\neq 0$. There will still be two competing ground states that depend on whether $\lambda_{3}$ or $\lambda_{4}$ (or both) is large or not. Due to the linear term (proportional to $\gamma_{1}$) in the ground state equations, the vanishing VEV from before is shifted slightly. For convenience we include $\gamma_{2}$ although it is allowed to vanish in the following ground state. In particular, we get
$\Phi_{\rm L}$ $=-\operatorname{sign}(\gamma_{1})\Phi_{\rm R}=w\mathbf{1}_{3},\quad w=\sqrt{\frac{m^{2}+2|\gamma_{1}|}{\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}+4\gamma_{2}}},$
$\langle V\rangle$ $=-\frac{3(m^{2}+2|\gamma_{1}|)^{2}}{2(\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}+4\gamma_{2})},$ (32)
when the condition
$2\left(1+\frac{2|\gamma_{1}|}{m^{2}}\right)^{2}>\left(1+\frac{\lambda_{3}+\lambda_{4}}{\lambda_{1}+\lambda_{2}}\right)\left(1+\frac{2\gamma_{1}^{2}(\lambda_{1}+\lambda_{2})}{m^{4}(\lambda_{3}+\lambda_{4}-\lambda_{1}-\lambda_{2}+4\gamma_{2})}\right),$ (33)
is satisfied and otherwise the following is the ground state
$\Phi_{\rm L}$ $=w_{\pm},\quad\Phi_{\rm R}=-\operatorname{sign}(\gamma_{1})w_{\mp},$
$w_{\pm}$ $=\frac{m}{\sqrt{\lambda_{1}+\lambda_{2}}}\sqrt{\frac{1}{2}\pm\epsilon\sqrt{1-\frac{16\gamma_{1}^{2}(\lambda_{1}+\lambda_{2})^{2}}{m^{4}(\lambda_{3}+\lambda_{4}-\lambda_{1}-\lambda_{2}+4\gamma_{2})^{2}}}},$
$\langle V\rangle$ $=-\frac{3m^{4}}{4(\lambda_{1}+\lambda_{2})}-\frac{6\gamma_{1}^{2}}{\lambda_{3}+\lambda_{4}-\lambda_{1}-\lambda_{2}+4\gamma_{2}},$ (34)
with the sign $\epsilon=\operatorname{sign}[\lambda_{3}+\lambda_{4}-\lambda_{1}-\lambda_{2}+4\gamma_{2}]$. The condition (33) choosing between the two types of ground states clearly gets complicated by the presence of $\gamma_{1}$ – the Josephson term, but the ground state structure is changed only little by $\gamma_{2}$. Interestingly, $\gamma_{2}$ does not affect the ground state condition when $\gamma_{1}=0$. One can check that setting $\gamma_{1}:=0$, the condition (33) reduces to the previous condition, i.e. $\lambda_{1}+\lambda_{2}>\lambda_{3}+\lambda_{4}$.
Clearly the ground state structure becomes only more complicated by turning on $\gamma_{3}\neq 0$. We will focus on the simplest case where we turn on $\gamma_{3}\neq 0$ but leave $\gamma_{1}=\gamma_{2}=0$. In this case, there is only a single ground state
$\Phi_{\rm L}$ $=-\Phi_{\rm R}=r\mathbf{1}_{3},\qquad r=\frac{1}{2}\sqrt{\frac{\lambda_{1234}-\sqrt{\lambda_{1234}^{2}-8m^{2}|\gamma_{3}|}}{|\gamma_{3}|}},$
$\langle V\rangle$ $=-\frac{\left(\lambda_{1234}-\sqrt{\lambda_{1234}^{2}-8m^{2}|\gamma_{3}|}\right)\left(16m^{2}|\gamma_{3}|-\lambda_{1234}\left(\lambda_{1234}-\sqrt{\lambda_{1234}^{2}-8m^{2}|\gamma_{3}|}\right)\right)}{32\gamma_{3}^{2}},$ (35)
provided that $\gamma_{3}$ is small enough:
$|\gamma_{3}|<\frac{\lambda_{1234}^{2}}{8m^{2}},$ (36)
and we have defined $\lambda_{1234}:=\sum_{i=1}^{4}\lambda_{i}$. If on the other hand, the above condition for $|\gamma_{3}|$ is not satisfied, the ground state becomes that of the partially unbroken phase:
$\left\{\begin{array}[]{l}\Phi_{\rm L}=v\mathbf{1}_{3}\\ \Phi_{\rm R}=\mathbf{0}_{3}\end{array}\right\}\ \ \textrm{or}\ \ \left\{\begin{array}[]{l}\Phi_{\rm L}=\mathbf{0}_{3}\\ \Phi_{\rm R}=v\mathbf{1}_{3}\end{array}\right\},\quad v=\frac{m}{\sqrt{\lambda_{1}+\lambda_{2}}},\quad\langle V\rangle=-\frac{3m^{4}}{4(\lambda_{1}+\lambda_{2})}.$ (41)
A comment in store about the $\gamma_{3}$ term, is that the potential theoretically has runaway directions that can be triggered for very large field values. This must however physically be just an artifact of the low-energy EFT.
3 Chiral vortex molecules
In this section, we numerically construct non-axisymmetric vortex configurations such as chiral vortex molecules and chiral vortices confined on a domain wall.
3.1 Vortex Ansätze for initial conditions
We seek solutions that describe vortex molecules that hence do not possess axial symmetry. Nevertheless, the initial configurations for our simulation need an Ansatz for each of the two vortices, which will be detailed here. The chiral $(1,0)$ vortex is given by
$\Phi_{\rm L}=v_{\rm L}\operatorname{diag}\left(f(r)e^{\mathrm{i}\theta},1,1\right),\qquad\Phi_{\rm R}=v_{\rm R}\operatorname{diag}(1,1,1),\qquad A=-\mathrm{i}\epsilon_{ij}\frac{x^{j}}{2r^{2}}a(r)T{\mathrm{d}}x^{i},$ (42)
with the matrix $T=\operatorname{diag}(\tfrac{2}{3},-\tfrac{1}{3},-\tfrac{1}{3})$ and the VEV of $\Phi_{\rm L}$ ($\Phi_{\rm R}$) being $v_{\rm L}$ ($v_{\rm R}$), depending on the ground state in question. Similarly, the chiral $(0,1)$ vortex is given by
$\Phi_{\rm L}=v_{\rm L}\operatorname{diag}(1,1,1),\qquad\Phi_{\rm R}=v_{\rm R}\operatorname{diag}\left(f(r)e^{\mathrm{i}\theta},1,1\right),\qquad A=-\mathrm{i}\epsilon_{ij}\frac{x^{j}}{2r^{2}}a(r)T{\mathrm{d}}x^{i},$ (43)
with the matrix $T=\operatorname{diag}(\tfrac{2}{3},-\tfrac{1}{3},-\tfrac{1}{3})$ and the VEV of $\Phi_{\rm L}$ ($\Phi_{\rm R}$) being $v_{\rm L}$ ($v_{\rm R}$), depending on the ground state in question.
In order to understand the normalization of the gauge field, we consider the split of the winding of the scalar field in eq. (42) into the global $\operatorname{U}(1)_{\rm B}$, $\operatorname{U}(1)_{\rm A}$, $\operatorname{SU}(3)_{\rm L}$ and $\operatorname{SU}(3)_{\rm R}$ with right action as well as the local $\operatorname{SU}(3)_{\rm C}$ with a left action:
$\operatorname{diag}\left(f(r)e^{\mathrm{i}\theta},1,1\right)$ $=e^{\mathrm{i}\beta\theta T}\operatorname{diag}\left(f(r),1,1\right)e^{\mathrm{i}\beta\theta T}e^{\mathrm{i}\alpha\theta}e^{\mathrm{i}\alpha\theta},$ (44)
$\operatorname{diag}\left(1,1,1\right)$ $=e^{\mathrm{i}\beta\theta T}\operatorname{diag}\left(1,1,1\right)e^{-\mathrm{i}\beta\theta T}e^{\mathrm{i}\alpha\theta}e^{-\mathrm{i}\alpha\theta}.$ (45)
The second equation, corresponding to the right field is trivially solved by any $\alpha$ and any $\beta$, whereas the first equation leads to two equations: $\frac{4}{3}\beta+2\alpha=1$ and $-\frac{2}{3}\beta+2\alpha=0$, yielding $\alpha=\frac{1}{6}$ and $\beta=\frac{1}{2}$. This means that the gauge part of the vortex carries the flux $\frac{1}{2}T$, which indeed is the normalization of the gauge field in eq. (42).
The boundary conditions for the two profile functions are
$f(0)=0,\qquad a(0)=0,\qquad f(\infty)=1,\qquad a(\infty)=1.$ (46)
A suitable initial guess for the profile functions takes the perturbative masses into account
$f_{\rm guess}(r)=\tanh(m_{\Phi}r),\qquad a_{\rm guess}(r)=\tanh(m_{A}r).$ (47)
For the initial state, we prepare a $(1,0)$ and a $(0,1)$ vortex with large enough separation that we can assume the following Abrikosov Ansatz
$\Phi_{\rm L}$ $=v_{\rm L}\operatorname{diag}\left(f(r_{\rm L})e^{\mathrm{i}\theta_{\rm L}},1,1\right),\qquad\Phi_{\rm R}=v_{\rm R}\operatorname{diag}\left(f(r_{\rm R})e^{\mathrm{i}\theta_{\rm R}},1,1\right),$
$A$ $=-\epsilon_{ij}\left(\frac{x_{\rm L}^{j}}{2r_{\rm L}^{2}}a(r_{\rm L})+\frac{x_{\rm R}^{j}}{2r_{\rm R}^{2}}a(r_{\rm R})\right)T{\mathrm{d}}x^{i},$ (48)
with $x+L+\mathrm{i}y=r_{\rm L}e^{\mathrm{i}\theta_{\rm L}}$ and $x-L+\mathrm{i}y=r_{\rm R}e^{\mathrm{i}\theta_{\rm R}}$ being two radial coordinates centered at the left and the right-hand side vortex, respectively.
The terms with the coefficients $\gamma_{1,2,3}$ give direct couplings between the left $\Phi_{\rm L}$ and right $\Phi_{\rm R}$ condensates. When all $\gamma_{1,2,3}$ are turned off, chiral vortices $(1,0)$ and $(0,1)$ are deconfined; they are attached by no domain walls. If we turn on at least one of $\gamma_{1,2,3}$, they are attached by one or two domain walls Eto:2021nle , as axion strings. This can be confirmed by substituting the Ansätze in eq. (42) or (43) into the potential term and by evaluating it at a large circle encircling a vortex. We then obtain a sine-Gordon model (when only one of $\gamma_{1,2,3}\neq 0$), a double sine-Gordon model (when $\gamma_{1,2}\neq 0$ and $\gamma_{3}=0$) and so on. The effective sine-Gordon models count the number of domain walls attached to the vortex that we are considering.555 In the two-Higgs doublet models as an extension of the Standard Model, a single non-Abelian vortex is attached by one or two domain walls depending on the parameters as shown by the same analysis Eto:2018hhg ; Eto:2018tnk . This model also admits a molecule of two non-Abelian vortex strings Eto:2021dca . In particular, in the case of $\gamma_{1,2}\neq 0$ and $\gamma_{3}=0$, the domain walls are non-Abelian sine-Gordon solitons carrying ${\mathbb{C}}P^{2}$ moduli Nitta:2014rxa ; Eto:2015uqa . [^6] The U$(N)$ non-Abelian sine-Gordon model appears also in the U$(N)$ chiral Lagrangian Eto:2021gyy and on a Josephson junction of two color superconductors Nitta:2015mma ; Nitta:2015mxa ; Nitta:2022ahj , and a sine-Gordon soliton can host SU$(N)$ Skyrmions as ${\mathbb{C}}P^{N-1}$ lumps Eto:2015uqa ; Eto:2023tuu ; Nitta:2022ahj . When a chiral vortex is attached to a non-Abelian sine-Gordon soliton their ${\mathbb{C}}P^{2}$ moduli match. The term with $\gamma_{3}\neq 0$ somehow “Abelianizes” the ${\mathbb{C}}P^{2}$ moduli.
3.2 Diagonal matrices
In ref. Eto:2021nle it was shown that $\mathbb{C}P^{2}$ moduli attract energetically. Under this assumption, choosing one vortex, say the left vortex in $\Phi_{\rm L}$ to be on diagonal form (without loss of generality), the right vortex in $\Phi_{\rm R}$ will align with the other and hence also be on diagonal form. Although we do not limit our simulations to diagonal matrices, a faster version that operates only using diagonal matrices can speed up the numerical investigations. For this reason we give the following Lemma.
Lemma 1
The variational equations (12)-(14) remain diagonal matrices when sourced by initial conditions that consist of diagonal matrices.
Proof: In order to facilitate the proof, we assume that the numerical method updates the fields by adding a constant times the equation of motion for that field to itself at every step. This is the case for the method given in sec. 3.3. It remains to check that the eom of eqs. (12)-(14) are diagonal matrices if the $\Phi_{\rm L}$, $\Phi_{\rm R}$, $A$ all are. Since the trace preserves the diagonal structure and the variational equations consist of products of diagonal matrices, only the symbols $\Xi_{{\rm L},{\rm R}}$ need to be checked. An explicit calculation reveals that $\Phi_{{\rm L},{\rm R}}$ being diagonal matrices reduces the variation of the determinant to
$(\Xi_{\rm L})_{\alpha a}=\frac{1}{2}\sum_{b,c=1}^{3}\epsilon_{\alpha bc}\epsilon_{abc}(\Phi_{\rm R})_{\alpha\alpha}(\Phi_{\rm L}^{\dagger}\Phi_{\rm R})_{bb}(\Phi_{\rm L}^{\dagger}\Phi_{\rm R})_{cc},$ (49)
which vanishes when $\alpha\neq a$ and similarly for $\Xi_{\rm R}$. $\square$
3.3 Numerical method
We will utilize the numerical method sometimes called arrested Newton flow, which similarly to relativistic dynamics accelerates towards the nearest local minimum of the energy functional. Unlike relativistic dynamics, the arrested part of the method is a continuous monitoring of the static energy (potential energy including field gradients) which sets the kinetic energy of the flow to zero once the energy increases. Specifically, we solve the equations
$\partial_{\tau}^{2}\Phi_{\rm L}=-\operatorname{eom}_{\Phi_{\rm L}},\qquad\partial_{\tau}^{2}\Phi_{\rm R}=-\operatorname{eom}_{\Phi_{\rm R}},\qquad\partial_{\tau}^{2}A=-\operatorname{eom}_{A},$ (50)
with an initial condition given in sec. 3.1 with $L$ typically set to $L:=4$. $\tau$ is not real time, but simply a parametrization of the flow. At every step of the flow, we compute the energy $E$ of eq. (3) and compare it with the energy of the previous step. If the energy has increased, we set $\partial_{\tau}\Phi_{\rm L}=\partial_{\tau}\Phi_{\rm R}=\partial_{\tau}A=0$.
The numerical computations are performed on square lattices with $1024^{2}$ lattice sites and the discrete derivatives are approximated using a 5-point stencil and a 4th-order numerical derivative. The lattice spacing is typically $0.0391$.
3.4 Numerical results
We will now perform numerical computations with the method described in sec. 3.3 and the initial conditions given in sec. 3.1. The main results are for the $(1,0)$ + $(0,1)$ vortices, each with winding in the 11 components of both the left and right scalar fields. Due to the large parameter space of the model (9-dimensional parameter space) and large number of possible initial conditions (relative $\mathbb{C}P^{2}$ coordinate), we cannot claim that we have exhausted all possibilities, but we get a general picture of how the set of chiral vortices behave. Also for simplicity, we will henceforth set $g:=1$ and $m:=\sqrt{2}$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: “Regular” (axisymmetric) non-Abelian vortex, i.e. no $\gamma$ terms turned on. In this figure $g=1$, $m=\sqrt{2}$, $\lambda_{1,2,3,4}=(2,0,0,0)$, $\gamma_{1,2,3}=(0,0,0)$. This vortex configuration has winding number $(1,1)$, the left and right fields’ vortices are coincident and the energy is axially symmetric (${\sf e}=0$).
Starting with the most simplistic choice of the potential parameters, we turn on only $\lambda_{1}$ (here and henceforth, the parameters $\lambda_{1}$ through $\lambda_{4}$ and $\gamma_{1}$ through $\gamma_{3}$ vanish when not turned on), i.e. we will set $\lambda_{1}:=2$. We are surprised to see that the two vortices attract and form a coincident vortex with total winding $(1,1)$, see fig. 1. The left and right fields’ vortices are completely coincident and the energy has eccentricity zero, where we define the eccentricity as [^7] We assume here that the configuration is elliptic with the major axis along the $x$-direction; otherwise the inverse of the fraction in the square root needs to be used.
${\sf e}=\sqrt{1-\frac{\int_{M}\star(x-x_{\rm CM})^{2}\mathcal{E}}{\int_{M}\star(y-y_{\rm CM})^{2}\mathcal{E}}},$ (51)
where center of mass is defined as $(x_{\rm CM},y_{\rm CM})=\int_{M}\star(x,y)\mathcal{E}$, the energy density is defined by $E=\int_{M}\star\mathcal{E}$ and $E$ is the energy given in eq. (3). Eccentricity zero corresponds to an axially symmetric energy configuration.
In the figure, we can clearly see the vortex “zero” in the 11 component of both the left and right scalar fields. Due to the gauge field being $\operatorname{SU}(3)$ and hence traceless, the gauge field must turn on the 22 and 33 components, which in turn induce nontrivial behavior in the 22 and 33 component of both the scalar fields. The “dip” in the 22 and 33 components of the scalar fields is, however, quite mild. The gauge field being $\operatorname{SU}(3)$ also implies that its field strength $F_{12}$ is traceless. The non-Abelian part of the field strength is, however, not gauge invariant. We thus display a gauge invariant quantity constructed out of the non-Abelian field strength as well as the two scalar fields
${\rm Im}\big{[}\operatorname{Tr}[F_{12}\Phi_{\rm L}\Phi_{\rm L}^{\dagger}]\big{]},\qquad{\rm Im}\big{[}\operatorname{Tr}[F_{12}\Phi_{\rm R}\Phi_{\rm R}^{\dagger}]\big{]},$ (52)
where taking the imaginary part is simply due to the convention of using anti-Hermitian gauge fields. The reason for multiplying by $\Phi_{\rm L}\Phi_{\rm L}^{\dagger}$ is that it is a matrix in color indices (i.e. with the flavors traced over).
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Confined chiral vortices. In this figure $g=1$, $m=\sqrt{2}$, $\lambda_{1,2,3,4}=(3,0,0,1)$, $\gamma_{1,2,3}=(-0.01,0,0)$. This vortex configuration has winding number $(1,1)$, the left and right fields’ vortices are coincident and the energy is axially symmetric (${\sf e}=0$).
In fig. 2, we turn on the $\lambda_{4}$ and also the Josephson term, i.e. $\gamma_{1}$. $\lambda_{4}$ does not have much impact on the vortex configuration, as long as it is smaller than $\lambda_{1}$. Taking it larger than $\lambda_{1}$ changes the ground state structure, but as we will see shortly, once it is of the same magnitude as $\lambda_{1}$, it will have an impact on the outcome.
By the logic that the Josephson term $(\gamma_{1})$ attaches one domain wall connecting the left and right chiral vortices, whereas the $\gamma_{2}$ term attaches two domain walls between the pair of vortices Eto:2021nle , we can predict that the tension of the Josephson wall will give rise to further attraction. In the example given in fig. 2, this is indeed the case, but since the vortices already want to be coincident, nothing much can be seen from turning on the Josephson term ($\gamma_{1}\neq 0$).
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Dipolar chiral vortices. In this figure $g=1$, $m=\sqrt{2}$, $\lambda_{1,2,3,4}=(1,0,0,1)$, $\gamma_{1,2,3}=(-0.01,0,0)$. This vortex configuration has winding number $(1,1)$, the left and right fields’ vortices form a dipole but the energy is axially symmetric (${\sf e}=0$). The scalar fields break the axial symmetry though.
In order to provoke some nontrivial structure out of the composite chiral vortices that like to be sitting in a coincident bound state, we leave the Josephson term on $\gamma_{1}\neq 0$, but lower $\lambda_{1}$ to the same level of $\lambda_{4}$, see fig. 3. This gives rise to a repulsion of the left and right chiral vortices, that are, however, still confined by the Josephson wall. Interestingly, although the scalar field clearly have distinct zeros (non-coincident) the energy is nevertheless axially symmetric. In that sense, just like a magnet, the bound state is dipolar with the two zeros forming the two poles of the state, which however remains axially symmetric in terms of the energy density. The eccentricity thus remains vanishing. In fig. 3 we also display $V_{1}$ which is the $\gamma_{1}$ part of $V$ (see eq. (4)).
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Chiral vortices with both Josephson and Josephson-squared terms turned on. In this figure $g=1$, $m=\sqrt{2}$, $\lambda_{1,2,3,4}=(3,0,0,1)$, $\gamma_{1,2,3}=(-0.01,0.01,0)$. This vortex configuration has winding number $(1,1)$, the left and right fields’ vortices form a dipole and the energy density is elliptic with eccentricity ${\sf e}=0.71$. The vortex dipolar molecule lives on an infinite Josephson wall, which however has vanishing energy density.
The way the Josephson term ($\gamma_{1}$) and its squared generalization ($\gamma_{2}$) connects to two vortices depends on the sign of the coefficients (or alternatively also on how the winding of the vortices is chosen; for example with $\gamma_{2}=\gamma_{3}=0$ the configurations are symmetric under $\gamma_{1}\to-\gamma_{1}$ and $\Phi_{\rm R}\to-\Phi_{R}$). In particular, the Josephson walls can form infinite domain walls. In fact, the non-axially symmetric vortex configuration shown in fig. 4 is created by having each vortex (i.e. left and right vortices) have a $\gamma_{1}$-wall that tends off to infinity, but meticulously choosing the $\gamma_{2}$ term with the exact same magnitude and opposite sign. This creates a small molecular (dipolar) bound state of left and right vortices, that are connected with one $\gamma_{1}$-wall, but have the other two $\gamma_{2}$-walls tending off to spatial infinity (recall that the $\gamma_{2}$ term attaches two walls to each vortex). Because of the fine tuned magnitude and opposite sign of the couplings $\gamma_{2}=-\gamma_{1}$, the energy of the wall tending off to infinity cancels out exactly (see $V_{1}$ and $V_{2}$ in fig. 4), leaving behind a confined albeit non-axially symmetric dipolar molecule of chiral vortices.888 This resembles a domain-wall bimeron in chiral magnets Amari:2024jxx .
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Chiral vortices as beats on a wall (Josephson double domain wall from the Josephson-squared term). In this figure $g=1$, $m=\sqrt{2}$, $\lambda_{1,2,3,4}=(3,0,0,1)$, $\gamma_{1,2,3}=(-0.01,0.05,0)$. This vortex configuration has winding number $(1,1)$, the left and right fields’ vortices form a dipole and the energy density is elliptic with eccentricity ${\sf e}=0.95$ (computed on the domain of the figure). The vortex dipolar molecule lives on an infinite Josephson wall with finite tension.
Now if we do not fine tune the cancellation between the Josephson term and its squared counterpart, we will create a domain wall with finite energy, see fig. 5. Increasing $\gamma_{2}$ creates an infinite double Josephson wall from each vortex. They are repelled slightly by the $\gamma_{1}$ term and confined by the $\gamma_{2}$ term, which however also gives an infinite contribution to the energy (if space is taken to be $\mathbb{R}^{2}$). This chiral vortex configuration is truly a molecule of $(1,0)$ and $(0,1)$ vortices with a visible separation (non-coincidence) in both the field zeros and in the energy density.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Confined vortices with the $\gamma_{3}$ term. In this figure $g=1$, $m=\sqrt{2}$, $\lambda_{1,2,3,4}=(3,0,0,1)$, $\gamma_{1,2,3}=(0,0,-0.01)$. This vortex configuration has winding number $(1,1)$, the left and right fields’ vortices are coincident and the energy is axially symmetric (${\sf e}=0$).
There are many other configurations that are very similar in nature to the examples we have selected above. As long as the ground state conditions allow for the same fully broken and symmetric (up to a phase or sign) ground state, many sets of coupling values give rise to very similar types of chiral vortex bound states. In fig. 6, we give an example of changing the $\gamma_{1}$ term for the $\gamma_{3}$ term in the configuration shown in fig. 2, yielding a very similar result. Another rule of thumb is that interpolating between $\lambda_{1}$ and $\lambda_{2}$ gives rise to the same type of configurations; the same holds true for interpolating between $\lambda_{3}$ and $\lambda_{4}$.
Figure 7: Energy during a flow for a configuration with the left vortex in the 11 component and the right vortex in the 22 component. The flow minimizes the energy, eventually aligning the two vortices in $\mathbb{C}P^{2}$. For details of the angle $\varphi$, see the text.
One aspect of the huge parameter space that we left out in the above results, is the fact that the left $(1,0)$ vortex and the right $(0,1)$ vortex can both be rotated around in $\mathbb{C}P^{2}$. The overall rotation (i.e. of both vortices) makes no difference in the energy density. However, there is a whole $\mathbb{C}P^{2}$ space of relative orientations, for example fixing the left vortex to be in the 11 component, the right vortex could be rotated to other points of $\mathbb{C}P^{2}$. We know from ref. Eto:2021nle that orthogonal vortices have larger energies than parallel ones. Let us fix the left vortex in the 11 component of $\Phi_{\rm L}$. The latter statement says that the energy is larger if we place the right vortex in the 22 component of $\Phi_{\rm R}$ than if we place it in the 11 component. One may contemplate what happens in between these two points of $\mathbb{C}P^{2}$. Using a rotation matrix of the form
$U(\varphi)=\begin{pmatrix}\cos\varphi&\sin\varphi&0\\ -\sin\varphi&\cos\varphi&0\\ 0&0&1\end{pmatrix},$ (53)
and rotating the fields as
$\Phi_{\rm L}$ $=v_{\rm L}\operatorname{diag}\big{(}f(r_{\rm L})e^{\mathrm{i}\theta_{\rm L}},1,1\big{)},\qquad\Phi_{\rm R}=v_{\rm R}U\operatorname{diag}\big{(}f(r_{\rm R})e^{\mathrm{i}\theta_{\rm R}},1,1\big{)}U^{\dagger},$
$A$ $=-\epsilon_{ij}\left(\frac{x_{\rm L}^{j}}{2r_{\rm L}^{2}}a(r_{\rm L})T+\frac{x_{\rm R}^{j}}{2r_{\rm R}^{2}}a(r_{\rm R})UTU^{\dagger}\right){\mathrm{d}}x^{i},$ (54)
with $x+L+\mathrm{i}y=r_{\rm L}e^{\mathrm{i}\theta_{\rm L}}$, $x-L+\mathrm{i}y=r_{\rm R}e^{\mathrm{i}\theta_{\rm R}}$ as usual and $f(r)$, $a(r)$ given in eq. (47), we can compute the energy as a function of $\varphi$. We choose a generic example with the couplings chosen as in fig. 2 and display the energy of the relative orientation rotated by $U$ of eq. (53) in fig. 7. The figure shows the energy during the flow towards the minimum of the energy functional while tracing the orientation $\varphi$, projected onto the $\operatorname{SO}(2)$ subspace of $\mathbb{C}P^{2}$. The angle $\varphi$ is not monotonic, but this is due to projecting onto the real subspace $\operatorname{SO}(2)$. Regardless of the projection chosen, the end result is that the vortices align their moduli to point in the same direction. Notice, however, that they point in the same direction, but are not pointing in the 11 direction due to the midpoint between 11 and 22 being some off-diagonal point.
4 Conclusion and outlook
In this paper, we have studied non-Abelian vortices in the CFL phase in dense 3-flavor QCD in a non-axially symmetric setting, especially searching for molecule-like or spatially nontrivial configurations with one vortex in the left condensation and one in the right condensation. The chiral vortices can be rotated around in $\mathbb{C}P^{2}$ and by studying the system with a left and a right vortex, this potentially leads to quite a large space of initial conditions. We have found, however, that the vortex moduli attract and the minimizers of the energy functional that we found were always in the parallel state, i.e. both the left and the right vortex pointing in the same direction inside $\mathbb{C}P^{2}$. Exploring the parameter space, we have found characteristic examples of bound states of left and right vortices, that break the axial symmetry and finally we have found the most molecular-like state of two chiral vortices on a chiral domain wall with two walls attached to each of the vortices – one of them binding the bound state and the other two ends tending off to infinity.
It would be interesting to get a better understanding of the vortex interactions in this model, which is quite complicated – especially at the nonlinear level, where the entire 9-dimensional parameter space kicks in. At the linear level, only the mass term and the Josephson term is probed by the scalar fields, but care must be taken for the non-Abelian gauge field to provide accurate predictions for the vortex interactions, which could probably be done by elaborating on the methods developed in ref. Speight:1996px .
Non-axisymmetric vortex configurations found in this paper break the axial symmetry spontaneously, yielding a rotational Nambu-Goldstone mode. Such a mode, called a twiston, can propagate along the vortex line, as a helium-3 superfluid. The transition between axisymmetric and non-axisymmetric configurations may affect thermal properties of a vortex lattice, which could be relevant for the core of neutron stars. Another direction is a possible construction of a non-Abelian vorton, i.e. a closed vortex string with a non-trivial twist.
In this paper, electromagnetic interactions are neglected for simplicity. They can be taken into account in the presence of a non-Abelian vortex, resulting in a nontrivial potential on the ${\mathbb{C}}P^{2}$ moduli Vinci:2012mc as well as an Aharanov-Bohm (AB) phase Chatterjee:2015lbf . The vortex molecules in this paper also ought to be studied in the presence of electromagnetic interactions. The scattering of charged particles such as electrons and charged CFL pions off of a vortex should exhibit nontrivial AB phases. In addition to the electromagnetic AB phase, single chiral non-Abelian vortices exhibit non-Abelian AB phases, i.e. when (quasi-)quarks encircle the vortex it picks up color nonsinglet AB phases Eto:2021nle , similarly to the case of non-Abelian Alice strings in the 2SC + dd phase Fujimoto:2020dsa ; Fujimoto:2021bes , which are also confined into a single non-Abelian vortex Fujimoto:2021wsr . The chiral non-Abelian vortex molecule may exhibit non-Abelian scattering if quarks can pass through between the two chiral vortices. One of the related nontrivial non-Abelian properties is the so-called topological obstruction, implying that generators of the unbroken symmetry in the ground state are not globally defined around the vortices Eto:2021nle . This might be relevant for the topological properties of the ground state(s).
Finally, beyond the GL effective theory, it can be shown in the Bogoliubov-de Gennes equation, describing (quasi-)quark degrees of freedom, that an axisymmetric non-Abelian vortex admits three Majorana fermion zeromodes in its core Yasui:2010yw ; Fujiwara:2011za ; Chatterjee:2016ykq . Such Majorana fermion zeromodes turn non-Abelian vortices into non-Abelian anyons Yasui:2010yh ; Hirono:2012ad ; Masaki:2023rtn .
Acknowledgments
S. B. G. thanks the Outstanding Talent Program of Henan University for partial support. The work of M. N. is supported in part by JSPS KAKENHI [Grant No. JP22H01221] and the WPI program “Sustainability with Knotted Chiral Meta Matter (SKCM2 )” at Hiroshima University (M. N.).